Find the center and radius of the circle given by the equation x^2 + y^2 - 8x + 2y - 11 = 0
Answer by nerdybill(5404)

x^2 + y^2 - 8x + 2y - 11 = 0
x^2 + y^2 - 8x + 2y = 11
Group x's and y's:
(x^2 - 8x) + (y^2 + 2y) = 11
Complete the square:
(x^2 - 8x + __ ) + (y^2 + 2y + __) = 11
(x^2 - 8x + 16 ) + (y^2 + 2y + 1) = 11 + 16 + 1
(x - 4 )^2 + (y + 1)^2 = 28
(x - 4 )^2 + (y - -1)^2 = 28
Center: (h,k) = (4, -1)
radius: sqrt(28) = 2sqrt(7)